From nch@roe.ac.uk Fri Feb 12 17:26:49 2010 Date: Fri, 12 Feb 2010 15:51:48 +0000 (GMT) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting, 12/02/2010 Minutes of WFAU VDFS Science Archive meeting: February 12th 2009 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, PMW, ETWS, NJC, RGM, RSC, RPB, MSH, MAR Apologies: JPE, AL, KTN DoNM: 10am, Friday February 19th 2010 in the VISTA Hut Actions discharged this week: ----------------------------- ACTION: RPB to check out availability/maturity of LTO4 drives and then liaise with PMW over purchase if required. Discharged; seems like a good time to get a drive to stick in the library unit; LTO3 will be deactivated, while the redundant LTO2 will be removed and kept on shelf. RPB and PMW will progress this. ACTION: ETWS to point RSC to an example command application that is failing in the list-driven photometry service Discharged Actions partly discharged but continuing: ----------------------------------------- ACTION: MAR, ETWS & NJC to put summary notes of the PSPIs meeting on the TWiki. NJC has done his bit; NCH suggested that once all contributions are in (asap please!) we need to distil into a prioritised list of probable/possible changes and enhancements, iterated with the PSPIs, and agreed amongst what remains of the developer effort. RGM noted that he has discussed with JPE a co-ordinating role when we've collated the change requests and prioritised. Actions carried forward from 05/02/10 meeting: ---------------------------------------------- None this week Specific points and new actions: -------------------------------- Project management: Nothing new to report this week. WFCAM & VISTA updates: NJC noted that we've had an enquiry from a member of the VSA-VIKING survey team concerning access of the dry-run data through the VSA, so are eagerly awaiting the v0.7 products from CASU. Comments and issues arising from CASU minutes: None this week. Networking: An SQL DB of the deep Stripe82 data has arrived from JHU courtesy of Alex Szalay, so the ops team will knit this into the archive and create crossmatches with UKIDSS DR7+ and VSA-VHS/VIKING etc. ACTION: RPB & NJC (with help when required from RSC) set up cross-neighbour tables between Stripe 82 and WSA/VSA surveys as appropriate. WSA/VSA Operations: ETWS noted: - Ingested metadata for 09B December. - Created JPGs for 09B December and re-created broken ones for 09B. - All tools for eyeballing checks for broken JPGs are in place so I will check through older semester as time allows. RPB noted: "The SDSS Stripe82 database has been received and copied onto our cluster servers. Will organise the necessary database entries to ensure that neighbour tables will be produced for DR7." MAR noted that DR7 QC1 is now in the cross-semester stage (so nearly done). The usual sanity checks will be applied to the merge logs and merged source lists before final curation stages prior to DR7 release. Hardware and Systems: MSH noted he's sent a spec for a new curation server to PMW who will look after the tendering. Software: NJC is continuing to generalise the non-survey auto-curation process and software for easier handling of likely schema spec-creep in the VSA PSs. NCH has finished analysing the relative astrometric performance of WFCAM for the production of proper motions for UKIDSS wide/shallow surveys using LAS and GPS DR6+, and asked RSC to work with him on the implementation in C++ to roll-out pm production in the standard curation codes (likely timescale DR8+ onwards). NJC asked that the implementation be as modular as possible to enable code re-use for the synoptic data model. ACTION: NCH and RSC to meet next week to sort out C++ implementation of proper motions in the merged source lists for UKIDSS Survey Data Release: As noted above, we're getting there with UKIDSS DR7... Non-survey Data Release: RPB reported: "Non Survey releases: In the current run the following non surveys were released: u/09a/d1 u/serv/1820 u/serv/1831 u/serv/1836 u/09a/2 There were problems with other releases, namely: u/serv/1802 u/serv/1801 u/09a/5 u/09a/1 u/09a/k1 but these will hopefully be released early next week [operational snafu, not problems with their data...] As requested, all u/cmp/1 data has now been copied into WFCAMOPENTIME." MAR is communicating with PIs for those non-surveys that are newly available. Astrogrid deployment & Data Analysis services: MSH noted that the firewall issues with VOSpace and Registry finally have been sorted out following tenacious negotiations with the network manager. Upgrades for TAP in the DSAs and security in the cone-search apps are in place and being tested - expecting deployment of new services imminently. Miscellaneous: RSC noted that Norman Gray is chairing a software session at NAM2010, (same afternoon as the Vista session, but a little before so we should go mob-handed to network and maybe make a presentation or two.) ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================